In the reaction: 2Na + 2H2O → 2NaOH+ H2

How many grams of hydrogen (H2


) are produced if 5.2 mol of Na are available?


1
Expert's answer
2021-05-20T07:54:36-0400

2Na + 2H2O → 2NaOH+ H2

According to the reaction, 2 moles of Na reacts to produce 1 mole of H2.

From here:

n(H2) = n(Na)/2 = 5.2 mol / 2 = 2.6 mol

Finally, the mass of H2 produced:

m(H2) = n(H2) × Mr(H2) = 2.6 mol × 2 g/mol = 5.2 g


Answer: 5.2 g